OpenStudy (anonymous):

plzzzzzzzz helpppppppp!!! 3+sqrt{z-9}=sqrt{z+6}

OpenStudy (anonymous):

Square both side and you get.. 9+z-9=z+6 z=z+6 its no solution.

OpenStudy (anonymous):

no it says their is a solution

OpenStudy (anonymous):

\[3+\sqrt{z-9}=\sqrt{z+6}\]

OpenStudy (anonymous):

did you write the problem correctly?

OpenStudy (anonymous):

ya thats how it looks